Jasmine deposits 1000 into a savings account paying nominal interest rate of 6 percent compounded three times a year. She decides to deposit the same amount at the beginning of each third of the year for 20 years. The interest earned from this account is withdrawn at the end of each year only and reinvested into another account earning AEIR 8%. Find the accumulated value of Jasmine’s money in both accounts at the end of the 20 years.
Answer #1
Every three deposits together with interest compounded
at 2%, can be treated as an annual deposit in the second
account.
Therefore, the annual year end deposit in the second
account = 1000*1.02*(1.02^3-1)/0.02 = $          3,121.61
It will be an ordinary annuity for 20 years, the FV of which = 3121.61*(1.08^20-1)/0.08 = $ 142,851.01
Answer: Accumulated value after 20 years = $142,851.01